OpenStudy (anonymous):

if the sales of a company increased exponentially by 50% from 1985 to 1990, by what percent will the sales increase from 1990 to 1996?

OpenStudy (anonymous):

any more information provided?

OpenStudy (anonymous):

no

OpenStudy (kainui):

Another business came around in 1993 and wiped out their profits so they went bankrupt and had no sales increase.

OpenStudy (kainui):

So let's consider the exponential equation. What does it look like? \[F(t)=Ae^{rt}\] So that tells us that A is our initial amount and F(t) is our final amount. Since we don't know the amounts but we do know the percents, you can divide out A to get a fraction, which translates to your percentage!

OpenStudy (anonymous):

can you show it? i cant figure this one out

OpenStudy (paxpolaris):

assuming they continue the same trend... you can PREDICT how much the sales will increase It increases by 1.5 times ... every 5 years .... IF you have a general formula for exponential function... you can plug these values in

OpenStudy (paxpolaris):

Using the formula above: ( i substituted \(\large c=e^r\)) \[\large F(t)= A \cdot c^t\]\[\implies \large 1.5 \cancel A= \cancel A \cdot c^5\]\[\implies \Large c=\sqrt[5]{1.5}\]

OpenStudy (paxpolaris):

So in 6 years ... it should increase by c^6 times : \[\Large = \left( \sqrt[5]{1.5} \right)^6=1.5^\left( 6/5 \right)\]

OpenStudy (paxpolaris):

62.67%